OpenStudy (keana):

need help quickly ! What is the sum of 7/x-1 + 9/x-8? a: 16/x-9 b: 16/x+8 c: 16x-64/x^2-9x+8; x ≠ -8, -1 d: 16x-64/x^2-9x+8; x ≠ 1, 8

OpenStudy (kropot72):

\[\frac{7}{x-1}+\frac{9}{x-8}=\frac{7(x-8)+9(x-1)}{(x-1)(x-8)}\] @keana Do you follow so far?

OpenStudy (keana):

@kropot72 i do follow yes !

OpenStudy (kropot72):

\[=\frac{7x-56+9x-9}{x ^{2}-9x+8}\] @keana Can you simplify the denominator?

OpenStudy (keana):

@kropot72 i don't think so ..

OpenStudy (kropot72):

Just group the like terms: 7x - 56 + 9x - 9 = 16x - 65. So the required sum becomes: \[\frac{16x-65}{x ^{2}-9x+8}\] x cannot equal 1 or 8, otherwise the denominator would become zero.

OpenStudy (kropot72):

@keana Sorry. I should have written "Can you simplify the numerator".

OpenStudy (keana):

@kropot72 i still would say no ..

OpenStudy (kropot72):

Grouping the like terms just means adding the 7x and 9x giving 16x, and combining the -56 and -9 giving -65

OpenStudy (keana):

@kropot72 alright ...

OpenStudy (kropot72):

Good! So the nearest answer is d, except the numerator shpuld be 16x - 65.

OpenStudy (keana):

@kropot72 oh ok thanks !

OpenStudy (kropot72):

You're welcome :)